# ALERT-Transformer: Bridging Asynchronous and Synchronous Machine Learning for Real-Time Event-based Spatio-Temporal Data

### v2

@@ 3 Experiments
Table 1: Classification performance and complexity on the DVS128Gesture dataset. Event representation meanings: (F=Frames, T=Tokens, E=Events). Online refers to the capacity of achieving the said accuracy in a continuous inference paradigm. Aync. refers to the event-driven processing nature of the network, where each event can be processed individually. TtA means Time to Accuracy.
@@
CNN + LSTM (Innocenti et al., 2021) F $500$ms+$t_{p}$ $97.7$% (SA)
$24$ms+$t_{p}$ $79.8$% (SA) [^3] The reported SA is obtained the model delivers a prediciton every 24ms, and latent vectors are reset every 20 samples.
@@ 3.1.1 Datasets
We validate our approach through two classification tasks: action recognition and binary classification. For action recognition, we utilize the 11-class DVS128Gesture dataset (Amir et al., 2017), consisting of 2 to 6-seconds recordings from $29$ users. For training (TE)LERT-Transformers, our input samples consist in Constant Count (CCIM) event streams of $Ne=8192$ events each, with the events being randomly sampled in sequence from the original recordings. A file usually contains far more than $8192$ events, thus ensuring a huge amount of possible random samples of $N_{e}$ consecutive events. For binary classification, the Prophesee’s N-Cars dataset (Sironi et al., 2018) is used. Each file is a $100$ms recording containing either a car or a scene without a car (“background”). In this case our input training samples correspond to one recording, using Constant Time (CTIM) event streams with variable number of events. We use the original train and test splits for both datasets, and during asynchronous inference simulation with ALERT, we treat the entire test files as continuous sequences of events.
@@ 3.1.2 Performance Metrics
Accuracy. It is common practice in the literature to process several samples sequentially before settling on a decision (a class) (Sabater et al., 2022; Wang et al., 2019a) – which we denote N-Voting Accuracy (NVA) –, or to estimate the class over full files (Peng et al., 2023; Innocenti et al., 2021). Because our model does not rely on any recurrence between two consecutive input samples, it can process each sample individually and independently. So, to compare with these works, we evaluate our model by voting over all predictions for each file, denoted File Voting Accuracy (FVA). However, to achieve ultra-low latency, one would need to be able to provide a decision with every single sample. Hence, we also show the average accuracy over all samples analyzed independently, which we denote as Sample Accuracy (SA). Using SA results in lower accuracy, but enables ultra-low latency.
@@ 3.2 Experimental Results
3.2.1 LERT-Transformer
Our RM model showcases noteworthy performance, with an FVA of $96.2$% on the DVSGesture dataset (see Table 1). When measured with SA the accuracy drops to 88.6%, but with much lower latency of $141.6$ms ($t_{in}$=$132$ms$\ +\ t_{p}$=$9.6$ms) on average. The RM is relatively complex, with the embedding running at $1.218$MFLOPs per event, and a total of $13.96$M parameters for the full (A)LERT-Transformer model (see Table 2). The LMM reduces complexity by $24$x to an impressive $0.566$M parameters, with the LERT module requiring $264$x less operation per event ($4.6$kFLOPs per event). It still shows a competitive FVA of $92.0$% on the action recognition task. On the binary classification task with the N-Cars dataset, the LMM achieves an accuracy of $85.6$% with $0.54$M parameters and an average of only $34.28$M total FLOPs per $100ms$.
3.2.2 ALERT-Transformer
Moving from the synchronous LERT to the TELERT module, we noticed a negligible relative accuracy drop in FVA of 0.4% and 0% for the RM and LMM, respectively (not shown in Table 1). This illustrates that time wrapping with sinusoidal encoding is an easy but efficient solution for bounded representation of time values. From synchronous LERT to ALERT (adding TE and decay) the accuracy relatively decreases by 2% and 3% for the RM and LMM versions. Nevertheless, this change in the model now allows for asynchronous updating of the tokens. The latency is entirely customizable as the tokens are continuously updated and can be processed on demand (the asynchronous to synchronous concept illustrated in Figure 1). The inference time $t_{p}$ for processing Ne=8192 events is $9.6$ms and $6.0$ms for the full RM and LMM, respectively. In the RM (LMM), (A)LERT takes $5.8$ms ($3.8$ms) against the $3.9$ms ($2.1$ms) of Transformer and Head. Because the (A)LERT module can be run event by event, we argue that using an event-driven asynchronous sparse AI accelerator could reduce the (A)LERT inference time to a few microseconds, thus enabling an on-demand ultra-low latency of $3.9$ms ($2.1$ms) for the RM (LMM).
@@ 3.2.3 Comparison with State-of-the-Art
Table 1 presents a comparative analysis with state-of-the-art for action recognition models. Many competitors lack the ability to process events asynchronously and cannot operate continuously (Online). PointNet++ (Wang et al., 2019a) still exhibits a low latency of $143$ms for an NVA accuracy of $95.3$% thanks to their sliding buffer voting and averaging strategy for improving accuracy. Note that our ALERT module could be deployed in their model, which would enable asynchronous PointNet++ proessing instead of the current update every 25ms. EventNet (Sekikawa et al., 2019), a comparable model with asynchronous to synchronous capability, has limited performance on complex tasks as it applies a single PointNet for the full input event stream spatial dimension. Accuracy measures on gesture recognition has thus not been reported by the authors. This was one of our motivation towards splitting the input window along the spatial dimension. (Sabater et al., 2022) propose a recurrent Transformer model that integrates events into frames and processes them every $24$ms. However, due to reccurence, predictions are originally obtained every 20 samples ($t_{in}=480$ms), introducing significant latency. When using their model by requesting a prediction every $24$ms (keeping the 20-sample recurrency) and applying SA the accuracy drops to $79.8$%33footnotemark: 3. In this situation, our ALERT-Transformer RM outperforms theirs with an accuracy of $84.6$%, but at the price of $1.218$MFLOPs/event. Even though our LMM requires only $0.0074$MFLOPs/event, in this case, it struggles with a $72.9$% SA. But the different hyperparameter searches realized during our study revealed a major cost/accuracy tradeoff (see Appendix B.1).
Table 2: Model complexity on the DVS128Gesture dataset.
Model #Param FLOPs $t_{p}$
(A)LERT Total (Ne=$2^{13}$)
RM $13.96$M $1.218$M/ev $9.42$G $9.6$ms
LLM $0.57$M $7.4$k/ev $0.06$G $6.0$ms
Also, when compared to state-of-the-art asynchronous models on the N-Cars binary classification task, our LMM showcases the lowest $0.0074$MFLOPs/event complexity while achieving a competitive accuracy (see Table 3). This represents a $75$% decrease in complexity compared to the previous state-of-the-art (Sironi et al., 2018), with a $5.11$% only decrease in accuracy.
3.3 Ablation Studies
@@
Figure 5: Influence of the patch size on accuracy and complexity (GFLOPs). Input resolution is $128\times 128$ pixels. The RM is set as the baseline, hence has a 1.0 relative accuracy.
Active Patch Threshold. Understanding the role of the activation threshold in the (A)LERT-Transformer is key for using it efficiently. This threshold – defined in number of events per patch – determines the portion of patches considered as active. Interestingly, according to Figure 6 comparable accuracies are achieved with both low and high thresholds. Nevertheless, higher thresholds result in less complex models because the FLOPs per processed event remains constant. In RM, the Transformer model is so small that the embedding part represents $94$% of the total FLOPs of the model for $Ne=8192$ events. Therefore, choosing a high threshold may be an important decision when a low complexity model with high accuracy is required.
@@
Input Sequence Length. The (TE)LERT module can be trained and evaluated in Constant Count or Constant Time input modes and is thus suited for processing input sequences of varying lengths ($Ne$). In Figure 7, we show how $Ne$ impacts the model performance. There seems to be an optimum point in between $Ne=8$k and $Ne=16$k events for the accuracy. However, the total complexity scales linearly with the number of processed events. Consequently, this parameter should be adjusted based on task-specific requirements. Note that the complexity of the ALERT module is not directly sensitive to the input sequence length used during training. However, the latter is indrectly related to the decay rate, which has an optimum value (see Appendix B for more details.).
@@ 3.4 Limitations and Potential Applications
While the outcomes of this study showcase the viability of the proposed ALERT module for event data in Transformer-based computer vision pipelines, there are some limitations and room for improvement. The current validation focuses on a Transformer architecture for classification tasks. However, the tokens do not have to be used only by Transformers, and our method could easily integrated with Convolutional Neural Networks (CNNs) or other architectures. Similarly, we are also working on the applicability of this learnt representation for other tasks than the ones used for benchmarking; such as generative solutions. We see several directions for exploiting our research: trying to improve accuracy by adopting advanced strategies akin to concurrent processing of multiple samples (Peng et al., 2023) or using of memory tokens (Sabater et al., 2022); or exploring PointNet-only architectures based on the ALERT module (Wang et al., 2019a; Sekikawa et al., 2019).
@@ A.2 Proof: Iterative Updating of Tokens
Here, $x_{i}$ represents the corresponding channel from each event feature. Now, if we want to do iterative updating of the tokens, we cannot recompute the $max$ over all values every time. Thus, we find a decay strategy where old channel values lose their relevance over time, instead of completely removing this values.
@@ B.3 Feature Generator Network Depth
Figure 11: Feature Generator’s Depth impact on the Model’s Performance. Overall,the depth of the (A)LERT module has a major influence on the accuracy of the network. The deeper the most accurate. But at the same time, network complexity increases. So a tradeoff needs to be made regarding the depth. On the other hand, the projection dimension does not have the same relationship with accuracy. Indeed, it seems that one should avoid using a too small projection width, but an optimum can be found. The same seems to be true for the expansion factor. The The two later points provide an optimism regarding the search for an optimal accuracy to cost ratio at a given depth. Note: the vertical red lines illustrate the RM configuration.
B.4 ALERT Inference Decay Rate
The decay rate is applied in the form of an exponential factor whenever a token channel has not been updated for a certain time. An optimal value exists for the decay rate, which should be tuned to simulate the behavior of the synchronous processing paradigm as truthfully as possible. Figure 12 shows the relationship between accuracy and decay rate for RM. Below the optimal point, our model forgets past information too quickly, above the optimal point too much past information is kept bringing confusion to the prediction. Note that this decay rate can also hint us toward the optimum length of the input event stream, as the decay rate emulates the duration of a sliding input window.

### v3

@@ 3 Experiments
Table 1: Classification performance and complexity on the DVS128Gesture dataset. Event representation meanings: (F=Frames, T=Tokens, E=Events). Online refers to the capacity of achieving the said accuracy in a continuous inference paradigm. Aync. refers to the event-driven processing nature of the network, where each event can be processed individually. TtA means Time to Accuracy, and it is the combination of input accumulation time ($\mathbf{t_{in}}$) and inference time ($\mathbf{t_{p}})$.
@@
EventTransAct (Blegiers et al., 2023) F $\rightarrow$ T $400$ms+$t_{p}$ $97.9$% (NVA)
$24$ms+$t_{p}$ $79.8$% (SA) [^3] The reported SA is obtained when the model makes a prediciton every 24ms, and latent vectors are reset every 20 samples.
CNN + LSTM (Innocenti et al., 2021) F $500$ms+$t_{p}$ $97.7$% (SA)
@@ 3.1.1 Datasets
We validate our approach through two classification tasks: action recognition and binary classification. For action recognition, we utilize the 11-class DVS128Gesture dataset (Amir et al., 2017), consisting of 2 to 6-seconds recordings from $29$ users. For training (TE)LERT-Transformers, our input samples consist in Constant Count (CCIM) event streams of $Ne=8192$ events each, with the events being randomly sampled in sequence from the original recordings. A file usually contains far more than $8192$ events, thus ensuring a huge amount of possible random samples of $Ne$ consecutive events. For binary classification, the Prophesee’s N-Cars dataset (Sironi et al., 2018) is used. Each file is a $100$ms recording containing either a car or a scene without a car (“background”). In this case our input training samples correspond to one recording, using Constant Time (CTIM) event streams with variable number of events. We use the original train and test splits for both datasets, and during asynchronous inference simulation with ALERT, we treat the entire test files as continuous sequences of events.
@@ 3.1.2 Performance Metrics
Accuracy. It is common practice in the literature to process several samples sequentially before settling on a decision (a class) (Sabater et al., 2022; Wang et al., 2019a) – which we denote N-Voting Accuracy (NVA) –, or to estimate the class over full files (Peng et al., 2023; Innocenti et al., 2021). Because our model does not rely on any recurrence between two consecutive input samples, it can process each sample individually and independently. So, to compare with these works, we evaluate our model by voting over all predictions for each file, denoted File Voting Accuracy (FVA). However, to achieve ultra-low latency, providing a decision with every single sample is crucial. Hence, we also show the average accuracy over all samples analyzed independently, which we denote as Sample Accuracy (SA). Using SA results in lower accuracy, but enables ultra-low latency.
@@ 3.2 Experimental Results
3.2.1 LERT-Transformer: Synchronous
Our RM model showcases noteworthy performance, with an FVA of $96.2$% on the DVSGesture dataset (see Table 1). When measured with SA, the accuracy drops to 88.6%, but with a much lower latency of $141.6$ms ($t_{in}$=$132$ms$\ +\ t_{p}$=$9.6$ms) on average. The RM is relatively complex, running at $1.299$MFLOPs per event, and a total of $13.96$M parameters for the full (A)LERT-Transformer model (see Table 2). The LERT module individually uses $1.218$MFLOPs/event and $1.41$M parameters. The LMM reduces this complexity by $24$x to an impressive $0.566$M parameters, with the LERT module requiring $264$x less operation per event ($4.0$kFLOPs per event), and $0.04$M parameters. The module together with the transformer runs at $7.4$kFLOPs per event. LMM still shows a competitive FVA of $92.0$% on the action recognition task. On the binary classification task with the N-Cars dataset, LMM achieves an accuracy of $85.6$% with $0.54$M parameters and an average of only $34.28$M total FLOPs per $100ms$. Note that the average FLOP count per sample varies across different datasets, as the number of events per sample differs depending on sensor resolution and scene dynamics. On the contrary, the FLOPs for processing a single event with our model always remain the same, independently of the dataset.
Table 2: Model complexity on the DVS128Gesture dataset. Here, LERT (RM/LMM) refers only to the LERT module part of the associated model, while RM and LMM refer to the entire model (LERT-Transformer). $t_{p}$ refers to the inference time, i.e. the time ALERT-Transformer takes to process a sample of $8192$ events.
Model #Param FLOPs per $t_{p}$
event sample
LERT (RM) $1.41$M $1.218$M $8.83$G $5.8$ms
RM $13.96$M $1.299$M $9.42$G $9.6$ms
LERT (LMM) $0.04$M $0.0040$M $0.03$G $3.9$ms
LMM $0.57$M $0.0074$M $0.06$G $6.0$ms
3.2.2 ALERT-Transformer: Asynchronous
Moving from the synchronous LERT to the TELERT module, we noticed a negligible relative accuracy drop in FVA of 0.4% and 0% for the RM ($95.8\%$) and LMM ($92.0\%$) respectively (not shown in Table 1). This illustrates that time wrapping with sinusoidal encoding is an easy but efficient solution for bounded representation of time values. From synchronous LERT to ALERT inferences (adding TE and decay) the FVA accuracy relatively decreases by 2% and 3% for the RM ($94.1\%$) and LMM ($89.2\%$) versions, respectively. Nevertheless, this change in the model now allows for asynchronous updating of the tokens. The latency is entirely customizable as the tokens are continuously updated and can be processed on demand (the asynchronous to synchronous concept illustrated in Figure 1). The inference times $t_{p}$ for processing $Ne=8192$ events are $9.6$ms and $6.0$ms for the full RM and LMM, respectively. In the RM (LMM), the (A)LERT module takes $5.8$ms ($3.9$ms) against the $3.8$ms ($2.1$ms) for Transformer and Head. Simulating this asynchronous low-latency scenario, the RM and LMM models achieve SA accuracies of $84.6\%$ and $72.9\%$, respectively. Because the (A)LERT module can be run event by event, we argue that using an event-driven asynchronous sparse AI accelerator could reduce the (A)LERT inference time to a few microseconds, thus enabling an on-demand ultra-low latency of $3.8$ms ($2.1$ms) for the RM (LMM).
@@ 3.2.3 Comparison with State-of-the-Art
Table 1 presents a comparative analysis with state-of-the-art for action recognition models. Many competitors lack the ability to process events asynchronously and cannot operate continuously (Online). PointNet++ (Wang et al., 2019a) still exhibits a low latency of $143$ms for an NVA accuracy of $95.3$% thanks to their sliding buffer voting and averaging strategy for improving accuracy. Note that our ALERT module could be deployed in their model, which would enable asynchronous PointNet++ processing instead of the current update every $25$ms. EventNet (Sekikawa et al., 2019), a comparable model with asynchronous to synchronous capability, has limited performance on complex tasks as it applies a single PointNet for the full input event stream spatial dimension. Accuracy measures on gesture recognition has thus not been reported by the authors. This was one of our motivations towards splitting the input window along the spatial dimension. EventTransAct (Blegiers et al., 2023) processes individual clips containing 16 frames of $5$ms. Every 5 clips, a prediction is obtained through mean aggregation, leading to a latency of $400$ms, and a NVA of $97.9$%. (Sabater et al., 2022) propose a recurrent Transformer model that integrates events into frames and processes them every $24$ms. However, due to reccurence, predictions are originally obtained every 20 samples ($t_{in}=480$ms), introducing significant latency. When using their model by requesting a prediction every $24$ms (keeping the 20-sample recurrency) and applying SA the accuracy drops to $79.8$%33footnotemark: 3. In this situation, our ALERT-Transformer RM outperforms theirs with an accuracy of $84.6$%, but at the price of $1.299$MFLOPs/event. Even though our LMM requires only $0.0074$MFLOPs/event, in this case, it struggles with a $72.9$% SA. But the different hyperparameter searches realized during our study revealed a major cost/accuracy tradeoff (see Appendix B.1).
Furthermore, when compared to state-of-the-art asynchronous models on the N-Cars binary classification task, our LMM showcases the lowest $0.0074$MFLOPs/event complexity while achieving a competitive accuracy (see Table 3). This represents a $75$% decrease in complexity compared to the previous state-of-the-art (Sironi et al., 2018), with only a $5.11$% decrease in accuracy.
3.3 Ablation Studies
@@
Figure 6: Influence of the patch size on accuracy ($\times$) and complexity ($+$) (GFLOPs). Input resolution is $128\times 128$ pixels. The RM is set as the baseline, hence has a 1.0 relative accuracy.
Active Patch Threshold. Understanding the role of the activation threshold in the (A)LERT-Transformer is key for using it efficiently. This threshold – defined in number of events per patch – determines the portion of patches considered as active. Interestingly, according to Figure 7 comparable accuracies are achieved with both low and high thresholds. Nevertheless, higher thresholds result in less complex models because the FLOPs per processed event remains constant, while less events are processed. In RM, the Transformer model is so small that the embedding part represents $94$% of the total FLOPs of the model for $Ne=8192$ events. Therefore, choosing a high threshold may be an important decision when a low complexity model with high accuracy is required.
@@
Input Sequence Length. The (TE)LERT module can be trained and evaluated in Constant Count or Constant Time input modes, and is thus suited for processing input sequences of varying lengths ($Ne$ or $\Delta_{t}$, respectively). In Figure 8a, we show how $Ne$ impacts the CCIM model’s performance. There seems to be an optimum point in between $Ne=8$k and $Ne=16$k events for the accuracy. However, the total complexity scales linearly with the number of processed events. Hence, our experiments were conducted using $Ne=8192$, which yields $96.2\%$ FVA. Similarly, for CTIM performance and complexity vary depending on $\Delta_{t}$ values (see Appendix B.4). The optimal point in this case is $\Delta_{t}=180$ms, where the model reaches $88.5\%$ SA and $96.5\%$ FVA. The average time span of a DVS Gesture input sample in CCIM with $Ne=8192$ fixed events, is of  $130$ms. It is worth noting that using $\Delta_{t}=130$ms in CTIM reaches $96.2\%$ FVA, proving the equivalence between both input modes. Consequently, the input mode and its length should be adjusted based on task-specific requirements.
Old Maximum Value Decay. The decay rate is applied in the form of an exponential factor whenever a token channel has not been updated for a certain time. An optimal value exists for this decay rate, which should be tuned to simulate the behavior of the synchronous processing paradigm as truthfully as possible. Figure 9 shows the relationship between accuracy and decay rate for RM. Below the optimal rate, our model forgets past information too quickly; above the optimal value too much past information is kept, bringing confusion to the prediction. In the limit where information is not forgotten, i.e. $\lambda=0$, the model reaches a sample accuracy of $25.9\%$, proving OMVD is a key component of the ALERT-Transformer.
Positional embeddings. The encodings of the patch coordinates, provide information about the patch tokens spatial distribution in space. Training the LERT-Transformer Reference Model (RM) on the DVSGesture dataset without these positional encodings, leads to accuracies of $80.0\%$ SA and $95.1\%$ FVA (versus $88.6\%$ SA, $96.2\%$ FVA with positional embeddings). During inference they can be built as a simple Look-Up table, introducing a negligible overhead to the total cost of the model.
@@ 3.4 Limitations and Potential Applications
While the outcomes of this study showcase the viability of the proposed ALERT module for event data in Transformer-based computer vision pipelines, there are some limitations and room for improvement. The current validation focuses on a Transformer architecture for classification tasks. However, the tokens do not have to be used only by Transformers, and our method could be easily integrated with Convolutional Neural Networks (CNNs) or other architectures. Similarly, we are also working on the applicability of this learnt representation for other tasks than those used for benchmarking, such as generative ones. We see several directions for exploiting our research: trying to improve accuracy by adopting advanced strategies akin to concurrent processing of multiple samples (Peng et al., 2023) or using of memory tokens (Sabater et al., 2022); or exploring PointNet-only architectures based on the ALERT module (Wang et al., 2019a; Sekikawa et al., 2019).
@@ 4 Conclusions
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work. We shortly mention two of them here specifically related to using an Event-based technology. Our work aims at bridging the gap between asynchronous, event-based signal processing and conventional signal processing. On one hand, in doing so, it facilitates the use of such low-latency system, for example for autonomous systems, and make them more robust and possibly safe (e.g. for obstacle detection). On the other hand, accelerating the use of event-based sensing and processing in computer vision might lead to systems requiring less power, which, at scale, is relevant for the environment. We do not foresee any negative impact of this work.
@@ A.2 Proof: Iterative Updating of Tokens
Here, $x_{i}$ represents the corresponding channel from each event feature. Now, if we want to do iterative updating of the tokens, we cannot recompute the $max$ over all values every time. Thus, we find a decay strategy where old channel values lose their relevance over time, instead of completely removing these values.
@@ B.3 Feature Generator Network Depth
Figure 13: Feature Generator’s Depth impact on the Model’s Performance. Overall,the depth of the (A)LERT module has a major influence on the accuracy of the network. The deeper the most accurate. But at the same time, network complexity increases. So a tradeoff needs to be made regarding the depth. On the other hand, the projection dimension does not have the same relationship with accuracy. Indeed, it seems that one should avoid using a too small projection width, but an optimum can be found. The same seems to be true for the expansion factor. The two later points provide an optimism regarding the search for an optimal accuracy to cost ratio at a given depth. Note: the vertical red lines illustrate the RM configuration.
B.4 Constant Time Input Mode
Figure 14: Duration of an Input Sample (in CTIM) impact on the Model’s Performance.